OpenStudy (anonymous):

How can I simplify : y = (x - 3)² - 1 Or is it already simplified?

OpenStudy (anonymous):

Apparently I need to combine like terms, but I don't see any :/

OpenStudy (anonymous):

y=x^2 - 9 -1 y=x^2 -10 ? i think that is how u do it

OpenStudy (anonymous):

Okay so that would be the simplified form?

OpenStudy (anonymous):

I just need to know for sure.

OpenStudy (anonymous):

yea cuz u need to distribute the exponent

jimthompson5910 (jim_thompson5910):

(x-3)^2 is not equal to x^2 - 9

OpenStudy (anonymous):

it isnt? oh i am wrong then

OpenStudy (anonymous):

@jim_thompson5910 So what would be the siplified form of this?

OpenStudy (anonymous):

Or is it already?

jimthompson5910 (jim_thompson5910):

you use the FOIL rule to expand out (x-3)^2 or (x-3)(x-3)

OpenStudy (anonymous):

it would be +9

jimthompson5910 (jim_thompson5910):

a table like this may help